OpenStudy (anonymous):

Find a so that the vectors v = i + aj and w = 3i + 2j are orthogonal.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

v and w are orthogonal if and only if the dot product of the two vectors is 0. v dot w = 0 (i + aj) dot (3i + 2j) = 0 1*3 + a*2 = 0 3 + 2a = 0 a = ???

OpenStudy (dan815):

@abb0t

OpenStudy (anonymous):

a = -3/2 ?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so if it were W = -4i -2j than it would be -4 -2a = 0 ?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

alright so it'd be -2, thank you.